# 6th Workshop on Patent Text Mining and Semantic Technologies (PatentSemTech) 2024

### v1

@@ 2 Methodology
Note that Equation 1 is a special case of Equation 4. The two equations are equivalent when there exists only a single positive pair with a score of 1, with all other pairs assigned a score of 0.
@@ 4 Experimental Results
where [OBJECT_NAME] represents the object’s description provided by DeepPatent2. These models provide significant improvements compared to the ViT and ResNet models from Table 1. This can potentially being attributed to the extensive and contextualized pretrained phases of CLIP. Additionally, language supervision further improves performance. Finally, we observe a similar performance trade-off between Patent ID and the higher hierarchical levels, as previously shown in Figure 1 for ViT Base and ViT Large. In the case of the CLIP models, the deterioration in performance at the Patent ID level is more pronounced, resulting from greater improvements in Subclass and Main Class levels.
Finally, the performance trade-off between CL and HMCL is more pronounced with CLIP, where our proposed HMCL method improves results at the Subclass and Main Class levels while showing reduced performance at the Patent ID level.
@@
Figure 3: First and second principal component of the embeddings of 5 different subclasses from the LIC system for the conventional Contrastive Learning (CL) and the Hierarchical Multi-Positive Contrastive Learning (HMCL) method. The Subclass 1042 and 1403 from LIC belong to the same Main Class, while the rest to different ones. The red boxes enclose the main regions of Subclass 1042 and 1403.

### v3

# 6th Workshop on Patent Text Mining and Semantic Technologies (PatentSemTech) 2025

@@ 2 Methodology
Note that Equation 1 is a special case of Equation 4. The two equations are equivalent when only a single positive pair exists with a score of 1, and all other pairs are assigned a score of 0.
@@ 4 Experimental Results
where [OBJECT_NAME] represents the object’s description provided by DeepPatent2. These models provide significant improvements compared to the ViT and ResNet models from Table 1. This can potentially being attributed to the extensive and contextualized pretrained phases of CLIP. Additionally, language supervision further improves performance. Finally, we observe a similar performance trade-off between Patent ID and the higher hierarchical levels, as previously shown in Table 1 for ViT Base and ViT Large. In the case of the CLIP models, the deterioration in performance at the Patent ID level is more pronounced, resulting from greater improvements in Subclass and Main Class levels.
@@
Figure 3: First and second principal component of the embeddings of 5 different subclasses from the LIC system for the conventional Contrastive Learning (CL) and the Hierarchical Multi-Positive Contrastive Learning (HMCL) method. The Subclass 1402 and 1403 from LIC belong to the same Main Class, while the rest to different ones. The red boxes enclose the main regions of Subclass 1402 and 1403.